OpenStudy (idealist10):

A space probe is to be launched from a space station 200 miles above Earth. Determine its escape velocity in miles/s. Take Earth's radius to be 3960 miles.

OpenStudy (idealist10):

@SithsAndGiggles

OpenStudy (anonymous):

If you consider the free body diagram of the probe, the only force acting on it (presumably) is the gravitational force of the Earth, so you have the DE \[m\frac{d^2x}{dt^2}=-\frac{mgR^2}{(x+R)^2}\] (where \(x(t)\) is the position of the probe, \(m\) is the mass of the probe, \(g\) is the acceleration due to gravity, and \(R\) is the radius of the Earth) To solve this you'll need two initial conditions (initial velocity and height), but you're only given one. According to this link, you can write the initial velocity as \(\sqrt{2gR}\). The derivation is all here: http://www.unc.edu/math/Faculty/mccombs/232/appdiffeqkey.pdf

OpenStudy (idealist10):

So after plugging the values, I got... \[\frac{ d ^{2}x }{dt ^{2}}=-\frac{ 9801}{338}\]

OpenStudy (idealist10):

Now what do I do?

OpenStudy (anonymous):

Your RHS is missing the dependent variable...

OpenStudy (idealist10):

So what should it be? What dependent variable did I miss?

OpenStudy (anonymous):

Plugging in all relevant values, the ODE is \[\frac{d^2x}{dt^2}=-\frac{3960^2g}{(x+3960)^2}\] (Since \(R\) is in units of miles and \(d^2x/dt^2\) is in miles per second per second, you need to convert \(g\) appropriately.)

OpenStudy (idealist10):

And then how do I solve for x(t) from there?

OpenStudy (anonymous):

According to this link ( http://www.unc.edu/math/Faculty/mccombs/232/appdiffeqkey.pdf ), the ODE is \[\frac{dv}{dt}=-\frac{3960^2g}{(x+3960)^2}\] I'm not sure if I interpreted the equivalence properly, i.e. if \(\dfrac{dv}{dt}=\dfrac{d^2x}{dt^2}\)...

OpenStudy (anonymous):

I think it holds, but it complicates the problem. Instead, you need to make a substitution so that the independent variable is \(x\), and not \(t\). To do this, notice that \(\dfrac{dv}{dt}=\dfrac{dv}{dx}\times\dfrac{dx}{dt}\). Since \(v=\dfrac{dx}{dt}\), you then have \(\dfrac{dv}{dt}=v\dfrac{dv}{dx}\). Now the equation is \[v\frac{dv}{dx}=-\frac{3960^2g}{(x+3960)^2}\] This equation is separable.

OpenStudy (idealist10):

But shouldn't we have a value for g?

OpenStudy (anonymous):

Yes, it's the constant \(9.81\dfrac{\text{m}}{\text{s}^2}\), Like I said earlier, you need to convert this to \(\dfrac{\text{mi}}{\text{s}^2}\) before you plug it in. If it helps, go ahead and solve \[v\frac{dv}{dx}=-\frac{gR^2}{(x+R)^2}\] first, then plug in the constants later.

OpenStudy (idealist10):

So I solved \[v \frac{ dv }{ dx }=-\frac{ 3960^{2} g}{ (x+3960)^{2} }\]

OpenStudy (idealist10):

and got\[\frac{ v^2 }{ 2 }=\frac{ 3960^2g }{ x+3960 }+C\]

OpenStudy (idealist10):

\[v^2=\frac{ 31363200g }{ x+3960 }+C\]

OpenStudy (anonymous):

Right. We don't have an explicit initial condition, so for the time being, we let the initial launch velocity be \(v_0\) when \(t=0\). At this time, though, we also have that \(x(0)=200\text{ mi}\). In other words, for \(v(x)\), you have \(v(200)=v_0\). Plugging this into the general solution, you get \[\frac{{v_0}^2}{2}=\frac{3960^2g}{200+3960}+C\] Solve for \(C\) in terms of \(v_0\).

OpenStudy (anonymous):

Once you find \(C\), you find the escape velocity by examining what happens to \(v_0\) as the height grows arbitrarily large; that is, compute \(\lim\limits_{h\to\infty}v_0\)

OpenStudy (anonymous):

(where \(h\) denotes the height; replace it with \(x\) for this context.)

OpenStudy (idealist10):

So I got\[C=v _{0}^2-\frac{ 31363200g }{ 4160 }\]

OpenStudy (anonymous):

Note that when solving for \(C\), it's no longer a general constant, so \(2C\not\to C\). Instead, the particular solution should be \[\frac{ v^2 }{ 2 }=\frac{R^2g }{ x+R}+\underbrace{\frac{{v_0}^2}{2}-\frac{R^2g}{200+R}}_{C}\] (I've replaced 3960 with \(R\) for a general radius). According to that link, when the height is maximized, the velocity is 0 (I'm still trying to figure out the physical interpretation of that, but I'll take their word for it at the moment), i.e. for max height \(x_{\text{max}}\), you have \(v=0\). Plugging this in gives \[0=\frac{R^2g}{x_{\text{max}}+R}+\frac{{v_0}^2}{2}-\frac{R^2g}{200+R}\] and solving for \(v_0\), you have \[v_0=\sqrt{2\left(\frac{gR^2}{200+R}-\frac{gR^2}{x_{\text{max}}+R}\right)}\] As \(x_{\text{max}}\to\infty\), you thus get \[v_0=\sqrt{\frac{2gR^2}{R+200}}\] where \(g=9.81\dfrac{\text{m}}{\text{s}^2}=0.006096\dfrac{\text{mi}}{\text{s}^2}\) and \(R=3960\text{ mi}\). In other words, the escape velocity should be about \(6.77\dfrac{\text{mi}}{\text{s}}\).

OpenStudy (idealist10):

Thank you!